### v2

# Efficient Algorithm for Sparse Fourier Transform of Generalized $q$-ary Functions

Abstract
Computing the Fourier transform of a $q$-ary function $f:\mathbb{Z}_{q}^{n}\rightarrow\mathbb{R}$, which maps $q$-ary sequences to real numbers, is an important problem in mathematics with wide-ranging applications in biology, signal processing, and machine learning. Previous studies have shown that, under the sparsity assumption, the Fourier transform can be computed efficiently using fast and sample-efficient algorithms. However, in most practical settings, the function is defined over a more general space—the space of generalized $q$-ary sequences $\mathbb{Z}_{q_{1}}\times\mathbb{Z}_{q_{2}}\times\cdots\times\mathbb{Z}_{q_{n}}$—where each $\mathbb{Z}_{q_{i}}$ corresponds to integers modulo $q_{i}$. Herein, we develop GFast, a coding theoretic algorithm that computes the $S$-sparse Fourier transform of $f$ with a sample complexity of $O(Sn)$, computational complexity of $O(Sn\log N)$, and a failure probability that approaches zero as $N=\prod_{i=1}^{n}q_{i}\rightarrow\infty$ with $S=N^{\delta}$ for some $10\leq\delta<$ 1. We show that a noise-robust version of GFast computes the transform with a sample complexity of $O(Sn^{2})$ and computational complexity of $O(Sn^{2}\log N)$ under the same high probability guarantees. Additionally, we demonstrate that GFast computes the sparse Fourier transform of generalized $q$-ary functions 8×8$\times$8 × faster using 16×16$\times$16 × fewer samples on synthetic experiments, and enables explaining real-world heart disease diagnosis and protein fitness models using up to 13×13$\times$13 × fewer samples compared to existing Fourier algorithms applied to the most efficient parameterization of the models as $q$-ary functions.
1 Introduction
Fourier analysis of $q$-ary functions, which map discrete sequences with $q$ alphabets to real numbers, plays a pivotal role in modeling and understanding complex systems across disciplines such as biology [1], signal processing [2], and machine learning [3]. Any $q$-ary function $f(\mathbf{m}):\mathbb{Z}_{q}^{n}\rightarrow\mathbb{R}$ can be expressed in terms of its Fourier transform $F[{\bf k}]$ as,
$f(\mathbf{m})=\sum_{\mathbf{k}\in\mathbb{Z}_{q}^{n}}F[\mathbf{k}]\omega^{\langle\mathbf{m},\mathbf{k}\rangle}\quad\mathbf{m}\in\mathbb{Z}_{q}^{n},$ (1)
where $\omega\coloneqq e^{\frac{2\pi j}{q}}$, and $\mathbf{m}$ and $\mathbf{k}$ are vectors in the $n$-dimensional ring of integers module $q$, denoted as $\mathbb{Z}_{q}^{n}=\{0,1,\dots,q-1\}^{n}$.
Computing the Fourier transform of a $q$-ary function $f$ is a challenging problem. Without any assumptions, this requires obtaining $N=q^{n}$ samples from $f$ and paying a computational cost of $O(N\log(N))$ to compute the transform using the seminal fast Fourier transform (FFT) algorithm [4]. This algorithm becomes computationally prohibitive for large $q$ or $n$. Fortunately, in practice, most $q$-ary functions have a sparse Fourier transform, meaning $F$ only has a few non-zero coefficients [5, 6]. The $S$-sparse Fourier transform of a $q$-ary function can be computed using efficient algorithms [7] with a sample complexity of $O(Sn)$, akin to the compressed sensing theory [8, 9, 10], and computational complexity of $O(Sn^{2}\log q)$, significantly faster than sparse regression methods, such as LASSO [11]. However, despite this progress, in most applications, $f$ is defined over alphabets of different sizes. For instance, consider a model that predicts patient outcomes using $n$ categorical features—such as age group, sex, or chest pain type. This model can only be represented as a function over the space $\mathbb{Z}_{q_{1}}\times\mathbb{Z}_{q_{2}}\times\cdots\times\mathbb{Z}_{q_{n}}$, where $q_{i}$ reflects the number of categories in feature $i$. Representing this model instead as a $q$-ary function is infeasible, since features often vary in their number of categories (grouping patients by age may require more categories than grouping by sex). As another example, consider a length-$n$ protein sequence, which can be naturally encoded as a $q$-ary sequence in $\mathbb{Z}_{q=20}^{n}$ for 20 amino acids. When predicting protein fitness from the sequence, it is common to assume only a fraction of the amino acids in each protein site are predictive [12]. In these cases, the function is more effectively modeled using a more generalized notion of $q$-ary function. We call such functions $f:\mathbb{Z}_{q_{1}}\times\mathbb{Z}_{q_{2}}\times\cdots\times\mathbb{Z}_{q_{n}}\rightarrow\mathbb{R}$, which map from the product of spaces with different alphabet sizes to real numbers, generalized $q$-ary functions.
Despite their importance, no efficient algorithm currently exists for computing the Fourier transform of generalized $q$-ary functions. This algorithm would enable several advances, including scalable methods to identify biological interactions in proteins [6, 13], explain machine learning models [14, 15], and regularize neural networks in the spectral domain [5, 3]. Existing workarounds naively require treating the underlying function as $q_{\text{max}}$-ary with $q_{\text{max}}=\max_{i}{q_{i}}$, which superficially inflates the problem’s dimensionality from $\prod_{i=1}^{n}q_{i}$ to $q_{\text{max}}^{n}$ with no change to the number of non-zero Fourier coefficients $S$. This leads to a substantial increase in sample and computational complexity. For example, modeling protein fitness using a generalized $q$-ary function by focusing on the subset of predictive amino acids predictive of a protein function results in more than twofold reduction in the problem dimensionality, from $q_{\text{max}}^{n=12}=8.9\times 10^{12}$ to $\prod_{i=1}^{n=12}q_{i}=8.6\times 10^{10}$ (see Fig. 4 in Section  7 of the Appendix). This effect is exacerbated as $n$ grows, making the computation of an otherwise feasible Fourier transform inaccessible due to sample and computational overheads.
Herein, we develop a novel algorithm to efficiently compute the sparse Fourier transform of a generalized $q$-ary function. We can express any generalized $q$-ary function $f$ as:
$f(\mathbf{m})=\sum_{\mathbf{k}\in\mathbb{Z}_{\mathbf{q}}}F[\mathbf{k}]\prod_{i=1}^{\smash{n}}\omega_{q_{i}}^{m_{i}k_{i}},\quad\mathbf{m}\in\mathbb{Z}_{\mathbf{q}},$ (2)
where $\mathbf{q}=[q_{1},q_{2},\dots,q_{n}]$, $\mathbb{Z}_{\mathbf{q}}=\mathbb{Z}_{q_{1}}\times\mathbb{Z}_{q_{2}}\times\cdots\times\mathbb{Z}_{q_{n}}$ for brevity, and $\omega_{q_{i}}=e^{\frac{2\pi j}{q_{i}}}$. In particular, with $q_{i}=q,\forall i$, Equations (1) and (2) become identical. This formulation naturally operates directly on alphabets of different sizes.
1.1 Contributions
We summarize the main contributions of our work:
  • We develop GFast, an efficient algorithm that computes the sparse Fourier transform of a generalized $q$-ary function with a sample complexity of $O(Sn)$ and computational complexity of $O(Sn\log N)$ in the noiseless setting. The GFast software is available on our GitHub repository [^1] https://github.com/amirgroup-codes/GFast
  • We develop a noise-robust version of GFast, dubbed NR-GFast, that computes the transform in the presence of additive Gaussian noise, with a sample complexity of $O(Sn^{2})$ and computational complexity of $O(Sn^{2}\log N)$.
  • We conduct large-scale synthetic experiments and demonstrate that GFast computes the sparse Fourier transform of generalized $q$-ary functions using up to 16×16$\times$16 × fewer samples and runs up to 8×8$\times$8 × faster than existing solutions.
  • We conduct real-world experiments to explain deep neural networks trained to predict heart disease diagnosis and protein fluorescence and demonstrate that NR-GFast computes the sparse Fourier transform using 13×13$\times$13 × fewer samples compared to existing algorithms.
1.2 Related Work
Computing the sparse Fourier transform of signals has a rich history, with much of the focus directed toward developing efficient algorithms for the discrete Fourier transform (DFT) [16, 17, 18, 19, 20, 21]. These algorithms leverage the principle of aliasing, where subsampling a signal in the time domain results in a linear mixing, or sketching, of the Fourier coefficients. Algorithms have been developed to efficiently sketch (encode) and recover (decode) Fourier coefficients. FFAST [16, 22] achieves this by inducing sparse graph alias codes in the DFT domain through a subsampling strategy guided by the Chinese Remainder Theorem and employing a belief-propagation decoder  [23, 24] to find the non-zero Fourier coefficients. Inspired by this coding-theoretic framework, algorithms are developed for the problem of computing the sparse Walsh Hadamard transform of pseudo-Boolean functions—a special case of a $q$-ary function with $q=2$ [25, 18, 26, 27]. More recently, the $q$-SFT algorithm [7] was developed to compute the sparse Fourier transform of a $q$-ary function for any integer $q$. GFast departs from all these previous works by directly subsampling in the $\mathbb{Z}_{\mathbf{q}}$ space. The resulting aliasing pattern induces a bipartite graph, which, unlike previous works, has an uneven degree distribution among the nodes (Sections 8.1 and 8.2 of the Appendix). Our theoretical results in GFast demonstrate that, surprisingly, the graph’s uneven degree distribution does not change the algorithm’s convergence guarantees in the asymptotic regime. GFast, in fact, computes the Fourier transform with a sample and time complexity that grows sublinearly in $N$, even in the presence of noise.
2 Problem Setup
Consider the generalized $q$-ary function $f({\bf m}):\mathbb{Z}_{\mathbf{q}}\rightarrow\mathbb{R}$ and let $N=\prod_{i=1}^{n}q_{i}$. Our goal is to find the Fourier transform of $f$ defined as,
$F[\mathbf{k}]=\frac{1}{N}\sum_{\mathbf{m}\in\mathbb{Z}_{\mathbf{q}}}f(\mathbf{m})\prod_{i=1}^{n}\omega_{q_{i}}^{-m_{i}k_{i}},\quad\mathbf{k}\in\mathbb{Z}_{\mathbf{q}},$ (3)
when $F[{\bf k}]$ is $S$-sparse. To establish theoretical guarantees, we make the following assumptions.
Assumption 1.
Let $\mathcal{S}\coloneqq\text{supp}(F)$ denote the support set of $F$ (i.e., the indices ${\bf k}$ of non-zero Fourier coefficients). We will make the following assumptions:
  1. 1.
    Each element in $\mathcal{S}$ is uniformly distributed and chosen randomly across $\mathbb{Z}_{\mathbf{q}}$.
  2. 2.
    The sparsity $S=|\mathcal{S}|$ is sub-linear in $N$, that is, $S=N^{\delta}$ for some $10\leq\delta<$ 1.
Assumption 2.
We also assume:
  1. 1.
    The Fourier coefficients $F[\mathbf{k}]$, for all $\mathbf{k}\in\mathcal{S}$, are sampled uniformly from the finite set $\mathcal{X}\coloneqq\{\rho,\rho\phi,\rho\phi^{2},\cdots,\rho\phi^{\kappa-1}\}$, where $\phi=e^{j\frac{2\pi}{\kappa}}$ for a constant phase offset $\kappa$ and strength factor $\rho$.
  2. 2.
    The signal-to-noise ratio (SNR) is written as: $\text{SNR}=\frac{||f||^{2}}{N\sigma^{2}}=\frac{||F||^{2}}{\sigma^{2}}=\frac{S\rho^{2}}{\sigma^{2}},$ and is assumed to be a constant.
  3. 3.
    We are given access to noisy samples $f(\mathbf{m})+w(\mathbf{m})$, where $w(\mathbf{m})\sim\mathcal{C}\mathcal{N}(0,\sigma^{2})$ is complex Gaussian noise.
3 GFast: Noiseless
We first analyze the case where samples drawn from $f$ are noiseless (i.e., $\sigma^{2}=0$). GFast finds the non-zero Fourier coefficients in three steps: 1) Subsampling and aliasing, 2) Bin detection, and 3) Peeling.
3.1 Sampling and Aliasing
GFast strategically aliases $f$ by computing small Fourier transforms using certain subsampling patterns. To do this, we create $C$ subsampling groups, each of size $b<n$, and take subsampled Fourier transforms over these samples. For each subsampling group $c$, we create the subsampling matrix $\mathbf{M}_{c}$ as an $n\times b$ matrix with the partial-identity matrix structure:
$\mathbf{M}_{c}=[\mathbf{0}_{b\times b(c-1)},\mathbf{I}_{b\times b},\mathbf{0}_{b\times(n-cb)}]^{T}.$ (4)
For each subsampling group, we denote the length-$b$ vector $\mathbf{b}_{c}$ as a size $b$ subvector of $\mathbf{q}$,
$\mathbf{b}_{c}\coloneqq\mathbf{M}_{c}^{T}\mathbf{q},$ (5)
and a set of $P$ offsets $\mathbf{d}_{c,p}\in\mathbb{Z}_{\mathbf{q}}$, where $p\in[P]$. For each $c,p$, the subsampled Fourier coefficients $U_{c,p}[\mathbf{j}]$ indexed by $\mathbf{j}\in\mathbb{Z}_{\mathbf{b}_{c}}$ are written as:
$U_{c,p}[\mathbf{j}]=\frac{1}{B_{c}}\sum_{\bm{\ell}\in\mathbb{Z}_{\mathbf{b}_{c}}}f(\mathbf{M}_{c}\bm{\ell}+\mathbf{d}_{c,p})\prod_{i=1}^{b}\omega_{(\mathbf{b}_{c})_{i}}^{-j_{i}\ell_{i}},\quad\mathbf{j}\in\mathbb{Z}_{\mathbf{b}_{c}},$ (6)
where $B_{c}=\prod_{i=1}^{b}(\mathbf{b}_{c})_{i}.$ In Section 8.1 of the Appendix, we show that this equals:
$U_{c,p}[\mathbf{j}]=\sum_{\mathbf{k}:\text{ }\mathbf{M}_{c}^{T}\mathbf{k}=\mathbf{j}}F[\mathbf{k}]\prod_{i=1}^{n}\omega_{q_{i}}^{(\mathbf{d}_{c,p})_{i}k_{i}},\quad\mathbf{k}\in\mathbb{Z}_{\mathbf{q}}.$ (7)
Let $\mathbf{s}_{c,\mathbf{k}}\in\mathbb{C}^{P}$ be the vector of offset signatures such that the $p^{\text{th}}$ value in $\mathbf{s}_{c,\mathbf{k}}$ is $\prod_{i=1}^{n}\omega_{q_{i}}^{(\mathbf{d}_{c,p})_{i}k_{i}}$. By grouping the subsampled Fourier coefficients as $\mathbf{U}_{c}[\mathbf{j}]=[U_{c,1}[\mathbf{j}],\ldots,U_{c,P}[\mathbf{j}]]$ with their respective offsets stacked as a matrix $\mathbf{D}_{c}\in\mathbb{Z}_{\mathbf{q}}^{P\times n}$, we rewrite Equation (7) as:
$\mathbf{U}_{c}[\mathbf{j}]=\sum_{\mathbf{k}:\text{ }\mathbf{M}_{c}^{T}\mathbf{k}=\mathbf{j}}F[\mathbf{k}]\mathbf{s}_{c,\mathbf{k}}.$ (8)
In the noiseless implementation, we set $P=n$ and $\mathbf{D}_{c}=\mathbf{I}_{n\times n}$.
Algorithm 1 GFast
1:$b,C,P,\{\mathbf{M}_{c}\}_{c\in[C]},\{\mathbf{D}_{c}\}_{c\in[C]}$
2:$\hat{F}\leftarrow\emptyset$
3:for each $c$ in $C$ do $\triangleright$ Subsampling
4:     for each $p$ in $P$ do
5:         $\mathbf{}U_{c,p}[\mathbf{j}]=\frac{1}{B_{c}}\sum_{\bm{\ell}\in\mathbb{Z}_{\mathbf{b}_{c}}}f(\mathbf{M}_{c}\bm{\ell}+\mathbf{d}_{c,p})\prod_{i=1}^{b}\omega_{(\mathbf{b}_{c})_{i}}^{-j_{i}\ell_{i}}$
6:     end for
7:end for
8:$\mathcal{S}=\{(c,\mathbf{j}):\text{Type}(\mathbf{U}_{c}[{\mathbf{j}}])=\mathcal{H}_{s}(\hat{\mathbf{k}},\hat{F}[\hat{\mathbf{k}}])\}$
9:$L=\{(c,\mathbf{j}):\text{Type}(\mathbf{U}_{c}[{\mathbf{j}}])=\mathcal{H}_{m}\}$
10:while $(|\mathcal{S}|>0)$ do $\triangleright$ Peeling
11:     for each $c$ in $C$ do
12:         for each $\mathbf{j}$ in $\mathbb{Z}_{\mathbf{b}_{c}}$ do
13:              if $\text{Type}(\mathbf{U}_{c}[\mathbf{j}])=\mathcal{H}_{S}(\hat{\mathbf{k}},\hat{F}[\hat{\mathbf{k}}])$ then
14:                  $\hat{F}\leftarrow\hat{F}\cup(\hat{\mathbf{k}},\hat{F}[\hat{\mathbf{k}})$
15:                  for each $c^{\prime}$ in $C$ do
16:                       $\mathbf{j}^{\prime}=\mathbf{M}_{c^{\prime}}^{T}\hat{\mathbf{k}}$
17:                       $\mathbf{U}_{c^{\prime}}[\mathbf{j^{\prime}}]\leftarrow\mathbf{U}_{c^{\prime}}[\mathbf{j^{\prime}}]-\hat{F}[\hat{\mathbf{k}}]\mathbf{s}_{c^{\prime},\hat{\mathbf{k}}}$
18:                       Update $\mathcal{S},L$
19:                  end for
20:              end if
21:         end for
22:     end for
23:end while
24:Return $\hat{F}$
3.2 Bin Detection and Recovery
After aliasing $f$, we use the subsampling groups to recover the original Fourier coefficients. Each $\mathbf{U}_{c}[\mathbf{j}]$ is a linear combination of Fourier coefficients $F[\mathbf{k}]$. The Fourier coefficients are recovered using a bin detection procedure, where the objective is to identify which $\mathbf{U}_{c}[\mathbf{j}]$ contains only one Fourier coefficient. To do this, in addition to the offset matrix $\mathbf{D}_{c}$, we choose a fixed delay $\mathbf{d}_{c,0}=\mathbf{0}_{n}$, where $\mathbf{0}_{n}$ is the vector of all 00’s of length $n$. This means that $U_{c,0}[\mathbf{j}]=\sum_{\mathbf{k}:\mathbf{M}_{c}^{T}\mathbf{k}=\mathbf{j}}F[\mathbf{k}]$. Using $U_{c,0}[\mathbf{j}]$, we are able to identify $\mathbf{U}_{c}[\mathbf{j}]$ as a zero-ton, singleton, or multi-ton bin based on the following criteria:
  • Zero-ton verification: $\mathbf{U}_{c}[\mathbf{j}]$ is a zero-ton (denoted by $\mathcal{H}_{Z}$) if $F[\mathbf{k}]=0$ such that $\mathbf{M}_{c}^{T}\mathbf{k}=\mathbf{j}$ (the summation condition in Equation (8)). In the presence of no noise, this is true when $U_{c,p}[\mathbf{j}]=0$ for all $p=1,\ldots,P$.
  • Singleton verification: $\mathbf{U}_{c}[\mathbf{j}]$ is a singleton (denoted by $\mathcal{H}_{S}$) if there exists only one $\mathbf{k}$ where $F[\mathbf{k}]\neq 0$ such that $\mathbf{M}_{c}^{T}\mathbf{k}=\mathbf{j}$. This implies that $|U_{c,1}[\mathbf{j}]|=|U_{c,2}[\mathbf{j}]|=\ldots=|U_{c,P}[\mathbf{j}]|$, meaning that in the noiseless case, $\mathbf{U}_{c}[\mathbf{j}]$ is a singleton if:
    $\left|\frac{U_{c,p}[\mathbf{j}]}{U_{c,0}[\mathbf{j}]}\right|=1,\quad p=1,2,\ldots,P.$
  • Multi-ton verification: $\mathbf{U}_{c}[\mathbf{j}]$ is a multi-ton (denoted by $\mathcal{H}_{M}$) if there exists more than one $\mathbf{k}$ where $F[\mathbf{k}]\neq 0$ such that $\mathbf{M}_{c}^{T}\mathbf{k}=\mathbf{j}$. In the noiseless case, this means that the following condition is met:
    $\left|\frac{U_{c,p}[\mathbf{j}]}{U_{c,0}[\mathbf{j}]}\right|\neq 1,\quad p=1,2,\ldots,P.$
With this, we define the function $\text{Type}(\mathbf{U}_{c}[\mathbf{j}])\rightarrow\{\mathcal{H}_{Z},\mathcal{H}_{M},\mathcal{H}_{S}(\mathbf{k},F[\mathbf{k}])\}$, which, given a bin, returns the type, along with a tuple of $(\mathbf{k},F[\mathbf{k}])$ if the bin is a singleton. If a singleton is found, $U_{c,0}[\mathbf{j}]=F[\mathbf{k}]$. This enables us to use the identity structure of $\mathbf{D}_{c}$ to solve for $\mathbf{k}$:
$\begin{bmatrix}\text{arg}_{q_{1}}[U_{c,1}[\mathbf{j}]/U_{c,0}[\mathbf{j}]]\\ \vdots\\ \text{arg}_{q_{n}}[U_{c,P}[\mathbf{j}]/U_{c,0}[\mathbf{j}]]\end{bmatrix}=\mathbf{D}_{c}\mathbf{k},$ (9)
where $\text{arg}_{q_{i}}\colon\mathbb{C}\rightarrow\mathbb{Z}_{q_{i}}$ is the $q_{i}$-quantization of the argument of a complex number defined as,
$\text{arg}_{q_{i}}(z):=\left\lfloor\frac{q_{i}}{2\pi}\text{arg}(ze^{\frac{j\pi}{q_{i}}})\right\rfloor.$ (10)
After obtaining the unknown vector $\mathbf{k}$, the value of $F[\mathbf{k}]$ is directly obtained as $F[\mathbf{k}]=U_{c,0}[\mathbf{j}]$.
3.3 Peeling Decoder
Upon determining the bin type, the final step uses a peeling decoder to identify more singleton bins. This problem is modeled with a bipartite graph, where each $F[\mathbf{k}]$ is a variable node, each $\mathbf{U}_{c}[\mathbf{j}]$ is a check node, and there is an edge between $F[\mathbf{k}]$ and $\mathbf{U}_{c}[\mathbf{j}]$ when $\mathbf{M}_{c}^{T}\mathbf{k}=\mathbf{j}$. The graph is a left-$C$ regular sparse bipartite graph with a total of $SC$ nodes. Fig. 1 shows an illustration of one such bipartite graph. The peeling algorithm is detailed in Algorithm 1. Upon identifying a singleton, for every subsampling group $c$ in $C$, we subtract $F[\mathbf{\mathbf{k}}]$ from every $\mathbf{U}_{c}[\mathbf{M}_{c}^{T}\mathbf{k}]$. This is equivalent to peeling an edge off the bipartite graph. After looping through all subsampled Fourier coefficients, this procedure is repeated until no more singletons can be peeled. Under Assumptions 1 and 2, we prove the following Theorem:
Theorem 1.
Given a function $f$ as Equation (2) that satisfies Assumptions 1 and 2, with each $B_{c}=O(S)$, $C=O(1)$, $P=n$, and $\mathbf{D}_{c}=\mathbf{I}_{n\times n}$, GFast recovers all Fourier coefficients in $\text{supp}(F)$ with probability at least $1-O(1/S)$ with a sample complexity of $O(Sn)$ and computational complexity of $O(Sn\log N)$ using subsampling matrices $\mathbf{M}_{c}$ in Equation (4).
Proof.
The proof is in Section 8.2 of the Appendix. ∎
4 GFast: Noise Robust
We develop a version of GFast that accounts for the presence of noise, where $\sigma^{2}>0$, dubbed NR-GFast. Instead of constraining $P=n$ and $\mathbf{D}_{c}=\mathbf{I}_{n\times n}$, given the hyperparameter $P_{1}$, we let $P=P_{1}n$ and $\mathbf{D}_{c}$ be a random matrix in $\mathbb{Z}_{\mathbf{q}}^{P\times n}$. For simplicity, we drop the subsampling group index $c$ and index $\mathbf{j}$ from $\mathbf{U}_{c}[\mathbf{j}]$ and write $\mathbf{U}=[\mathbf{U}_{1},\cdots\mathbf{U}_{p}]^{T}$. For each group, let $\mathbf{S}=[\cdots\mathbf{s_{k}}\cdots]\in\mathbb{C}^{P\times N}$ be the matrix of offset signatures. In the presence of noise, $\mathbf{U}=\mathbf{S}\bm{\alpha}+\mathbf{W}$, where $\bm{\alpha}[\mathbf{k}]=F[\mathbf{k}]$ if $\mathbf{M}^{T}{\mathbf{k}}=\mathbf{j}$, otherwise $\bm{\alpha}[\mathbf{k}]=0$. $\mathbf{W}$ is a matrix of complex multivariate Gaussian noise with zero mean and covariance $\nu_{c}^{2}\mathbf{I}$, such that $\nu_{c}^{2}=\sigma^{2}/B_{c}$.
NR-GFast uses random offsets that individually recover each $\mathbf{k}$ index based on repetition coding. Given a set of $P_{1}$ random offsets $\mathbf{d}_{p\in[P_{1}]}$ such that $P=P_{1}n$, we modulate each offset with each column of the identity matrix such that we have $n$ offsets $\mathbf{d}_{p,r}:$
$\mathbf{d}_{p,r}=\mathbf{d}_{p}\oplus_{q}\mathbf{e}_{r},\forall p\in[P_{1}],\quad\forall r\in n.$ (11)
This generates $C$ matrices $\mathbf{D}_{c}\in\mathbb{Z}_{\textbf{q}}^{P\times n}$. We then use a majority test to estimate the $r^{\text{th}}$ value of $\hat{\mathbf{k}}$, where $\hat{\mathbf{k}}$ is the estimated index of $\mathbf{k}$:
$\hat{k}_{r}=\operatorname*{arg\,max}_{a\in\mathbb{Z}_{q_{r}}}\sum_{p\in[P_{1}]}\mathbbm{1}\{a=\text{arg}_{q_{r}}[U_{p,r}/U_{p}]\},$ (12)
where $\mathbbm{1}$ denotes the indicator function, and $\hat{k}_{r}$ denotes the value $a\leq q_{r}$ that is the maximum likelihood estimate of $q_{r}$.
Figure 1: An example of a bipartite graph for computing the sparse Fourier transform of a generalized $q$-ary function with $n=4$, $b=2$, and $\mathbf{q}=[3,3,3,2]$, where $\mathbf{b}_{1}=[3,3]$ (blue squares) and $\mathbf{b}_{2}=[3,2]$ (red squares). Dashed lines represent edges that are peeled in the first iteration of decoding in GFast. Existing algorithms for $q$-ary functions use the $\mathbf{b}_{2}=[3,3]$ (red and gray squares) grouping, which inflates the dimension.
Then, for some $\gamma\in(0,1)$, bin detection is modified using the following criteria:
  • Zero-ton verification: $\text{Type}(U_{c}[j])=\mathcal{H}_{Z}\frac{1}{P}\|\mathbf{U}_{c}[\mathbf{j}]\|^{2}\leq(1+\gamma)\nu_{c}^{2}$.
  • Singleton verification: After ruling out zero-tons, we estimate $(\hat{\mathbf{k}},\hat{F}[\hat{\mathbf{k}}])$, where $\hat{F}[\hat{\mathbf{k}}]$ is the estimated Fourier coefficient corresponding to $\hat{\mathbf{k}}$, using the procedure described in the noiseless case. We verify if $\mathbf{U}_{c}[\mathbf{j}]$ is a singleton if $\frac{1}{P}\|\mathbf{U}_{c}[\mathbf{j}]-\hat{F}[\hat{\mathbf{k}}]\mathbf{s}_{c,\hat{\mathbf{k}}}\|^{2}\leq(1+\gamma)\nu_{c}^{2}$.
  • Multi-ton verification: If neither $\mathrm{Type}(U_{c}[j])=\mathcal{H}_{Z}$ nor $\text{Type}(\mathbf{U}_{c}[\mathbf{j}])=\mathcal{H}_{S}$, then $\text{Type}(\mathbf{U}_{c}[\mathbf{j}])=\mathcal{H}_{M}$.
Using the above setup, the random offsets satisfy the following proposition:
Proposition 1.
Given a singleton bin $(\mathbf{k},F[\mathbf{k}])$, the $q_{r}$-quantized ratio as defined in Equation (10) for
$U_{p}$ $=F[\mathbf{k}]\prod_{i=1}^{n}\omega_{q_{i}}^{(\mathbf{d}_{p})_{i}k_{i}}+W_{p},$ (13)
$U_{p,r}$ $=F[\mathbf{k}]\prod_{i=1}^{n}\omega_{q_{i}}^{(\mathbf{d}_{p,r})_{i}k_{i}}+W_{p,r},$
satisfies $\text{arg}_{q_{r}}[U_{p,r}/U_{p}]=\langle\mathbf{e}_{r},\mathbf{k}\rangle\otimes_{q_{r}}Z_{p,r}$, where $Z_{p,r}$ is a random variable over $\mathbb{Z}_{q_{r}}$ with a probability $p_{i}\coloneqq\mathbb{P}(Z_{p,r}=i)$. Let $\sum_{i\neq 0}p_{i}\coloneqq\mathbb{P}_{e}$. $\mathbb{P}_{e}$ is upper bounded by $2e^{-\frac{\zeta}{2}\text{SNR}}$ for $\zeta=\eta_{c}\sin^{2}(\pi/2q_{r})$.
Proof.
The proof is in Section 8.3 of the Appendix. ∎
Using Proposition 1, we prove the following Theorem:
Theorem 2.
Given a function $f$ as Equation (2) that satisfies Assumptions 1 and 2, with each $B_{c}=O(S)$, $C=O(1)$, $P=n^{2}$, and $\mathbf{D}_{c}$ defined as in Equation (11), NR-GFast recovers all Fourier coefficients in $\text{supp}(F)$ with probability at least $1-O(1/S)$ with a sample complexity of $O(Sn^{2})$ and computational complexity of $O(Sn^{2}\log N)$ using subsampling matrices $\mathbf{M}_{c}$ in Equation (4) as $n$ increases and if each $q_{i}=O(1)>1$.
Proof.
The proof is in Section 8.4 of the Appendix. ∎
5 Experiments
We first evaluate the performance of GFast on synthetic data by considering the scenario where four different generalized $q$-ary functions taking in an $n=20$ length sequence have $q_{\text{max}}=7$ (the distribution of alphabets is shown in the legend of Fig. 2a). We run $q$-SFT on one function by setting $q_{i}=q_{\text{max}},\forall i\in 1\dots n$. Following Assumptions 1 and 2, we synthetically generate an $S$-sparse Fourier transform $F$, where $\mathcal{S}=\text{supp}(F)$ is chosen uniformly at random in $\mathbb{Z}_{q_{1}=7}\times\mathbb{Z}_{q_{2}=3}\times\mathbb{Z}_{q_{3}=3}\times\mathbb{Z}_{q_{4}=3}\times\mathbb{Z}_{q_{5}=4}\times\cdots\times\mathbb{Z}_{q_{20}=4}$ with values according to $F[\mathbf{k}]=\rho e^{-j\Omega[\mathbf{k}]}\text{ if }\mathbf{k}\in\mathcal{S},\text{ else }0$, where $\Omega[\mathbf{k}]$ are independent and random variables sampled from $[0,2\pi)$. We evaluate performance using normalized mean-squared error (NMSE), which is calculated as: $\text{NMSE}=\frac{||\hat{F}-F||^{2}}{||F||^{2}}=\frac{||\hat{f}-f||^{2}}{||f||^{2}}$, where $\hat{f}$ is the estimated function using the recovered Fourier coefficients. In all experiments, we set $C=\lfloor n/b\rfloor$. Fig. 2a and 2b compare GFast and NR-GFast with the noiseless and noise-robust versions of $q$-SFT in terms of sample complexity, respectively. We utilize $q$-SFT’s partial-identity subsampling matrix implementation to match the aliasing pattern derived in Equation (7). In the noiseless setting, $\rho=1$, $S=1000$, and $\sigma^{2}$ such that $\text{SNR}=100$dB. To vary the sampling patterns of $\mathbf{q}$ in the noiseless implementation of GFast, we additionally permute $\mathbf{q}$ and each vector of $\mathbf{k}$ five times per experimental instance. For both algorithms, we vary $b$ from 1 to 5 to get different sample counts. In the noisy setting, we set $\rho=1$, $S=1000$, and $\sigma^{2}$ such that $\text{SNR}=10$dB. We vary $b$ from 1 to 5 and $P_{1}$ from 18 to 20 in both algorithms. Each $q$-SFT and GFast experimental instance is run over three random seeds. When GFast achieves an NMSE of less than $0.01$, there is an average difference of $0.96\pm 0.06$ and $0.92\pm 0.06$ in NMSE at the closest sample complexity of $q$-SFT in the noiseless and noisy settings, respectively. When both GFast and $q$-SFT achieve an NMSE of less than $0.01$, GFast uses up to 16×16$\times$16 × fewer samples and runs up to 8×8$\times$8 × faster.
Figure 2: NMSE of GFast compared to $q$-SFT on synthetically generated data, where the Fourier coefficients live in $\mathbb{Z}_{q_{1}=7}\times\mathbb{Z}_{q_{2}=3}\times\mathbb{Z}_{q_{3}=3}\times\mathbb{Z}_{q_{4}=3}\times\mathbb{Z}_{q_{5}=4}\times\cdots\times\mathbb{Z}_{q_{20}=4}$, in the (a) noiseless and (b) noisy settings at various sample complexities. When achieving an NMSE of less than $0.01$, GFast uses up to 16×16$\times$16 × fewer samples and runs up to 8×8$\times$8 × faster than $q$-SFT.
We additionally show the performance of NR-GFast in learning a real-world function. We train a multilayer perceptron $f$ to predict the presence of heart disease using data from [28]. $f$ takes $n=12$ categorical features as an input. Fig. 3a shows the number of categories per feature, where the maximum alphabet size is $q=5$. Treating this problem as a $q$-ary function with $q=5$ artifically inflates the problem’s dimensionality from $\prod_{i=1}^{n=12}q_{i}=4.5\times 10^{6}$ to $q^{n}=5^{12}\sim 2.4\times 10^{8}$. Moreover, evaluating $f$ on out-of-domain inputs $\mathbf{x}\in\mathbb{Z}_{q}^{n}\setminus\mathbb{Z}_{\mathbf{q}}$ unnecessarily increases sample complexity, since computing $f(\mathbf{x})$ requires freezing the input features $\mathcal{V}=\left\{i\in\{1,\dots,n\}\mid x_{i}\in\mathbb{Z}_{q_{i}}]\right\}$ that exist in $\mathbb{Z}_{\mathbf{q}}$, and taking an expectation over the remaining samples: $f(\mathbf{x})=\mathbb{E}_{\mathbf{m}\sim\mathrm{Unif}(\mathbb{Z}_{\mathbf{q}})}\left[f(\mathbf{m})\,\middle|\,\mathbf{m}_{\mathcal{V}}=\mathbf{x}_{\mathcal{V}}\right]$. We run NR-GFast over inputs in $\mathbb{Z}_{\mathbf{q}}$ and $q$-SFT using $q=5$ and $n=12$, and calculate the test NMSE of both algorithms on $10,000$ samples chosen uniformly and randomly from $\mathbb{Z}_{\mathbf{q}}$. We vary $b$ from 1 to 3 and $P_{1}$ from $3$ to $12$ in both algorithms to get different sample complexities.
Fig. 3b compares NR-GFast with $q$-SFT. Despite NR-GFast taking 13×13$\times$13 × fewer samples than $q$-SFT, NR-GFast achieves a minimal NMSE of $0.37$ compared to $q$-SFT achieving a minimal NMSE of $0.42$. Given a sample budget of $10^{5}$ samples, $q$-SFT achieves a significantly worse minimal NMSE ($0.53$) than NR-GFast. On an MLP trained to predict protein floresence (see Fig. 4 in Section 7 in the Appendix), NR-GFast again outperforms $q$-SFT given a sample budget (with an NMSE of $0.42$ and $0.54$, respectively). Together, these results highlight the sample efficiency of NR-GFast in computing the sparse Fourier transform of generalized $q$-ary functions.
Figure 3: (a) Distribution of categories per feature that comprise the generalized $q$-ary function in NR-GFast for $n=12$ compared to the $q$-ary function where $q=5$. (b) Performance of NR-GFast and $q$-SFT on the presence of heart disease as predicted by a trained MLP. NR-GFast uses up to 13×13$\times$13 × fewer samples than $q$-SFT.
6 Conclusion and Future Work
In this paper, we introduced GFast, a fast and efficient algorithm for computing the sparse Fourier transform of generalized $q$-ary functions. Our theoretical analysis and empirical experiments demonstrate that GFast outperforms existing algorithms limited to $q$-ary functions by achieving faster computation and requiring fewer samples, especially when the alphabet size distribution is non-uniform.
Several promising research directions remain to be explored. For example, our theoretical analysis with noise assumes bounded $q_{i}$, and our experiments are limited to alphabet sizes up to $12$. A significant increase in $q$, for example, to several thousand in language processing applications, will degrade GFast’s performance. Developing algorithms that are robust to large alphabets is an important future challenge. Additionally, leveraging GFast to develop methods for explaining large-scale machine learning models of generalized $q$-ary functions—such as those used in biology and natural language processing—offers another exciting avenue for further research.